# Predicate Invention from Pixels via Pretrained Vision-Language Models

### v1

@@ 5 Experiments
    ViLa: A baseline that does not learn abstractions but uses a VLM to plan on test tasks directly (Hu et al. 2023a). We test two variants of this: ViLa-pure which is the original approach as implemented by Hu et al. (2023a) adapted to our setting, and ViLa-fewshot, which additionally provides the demonstrations as part of the few-shot prompt to give the VLM some context of what kinds of plans it should return for testing tasks.
Experimental Setup. Each seed varies the number and locations of objects in the train and test tasks, and we average results over $5$ random seeds for each task distribution. For each seed, we sample 10 test tasks. We use GPT-4o (OpenAI 2023) as the VLM for all approaches. For our tasks Burger, as mentioned in section 4.3, we early stop the predicate subselection optimization when the hill-climbing objective value goes below $2000$ (a value we found by running pix2pred on a validation set). We evaluate each approach by providing the test task and allowing the approach to make and execute exactly one single plan.
Results and Analysis. Figure 7 shows our success rate plots. pix2pred outperforms previous approaches on 3 out of 4 of the task distributions by a wide margin. In particular, pix2pred seems to generalize to more complex tasks much better than using few-shot learning with a VLM (i.e., ViLa) (Q1). In the short-horizon kitchen task that requires 2 actions, we find that ViLa with demonstrations provided performs comparably to pix2pred. However, the performance of this baseline drops-off significantly in the longer-horizon burger tasks that require more than 15 actions. In particular, we found that ViLa seems to solve testing tasks by pattern-matching aspects of the training demonstrations and is unable to solve test tasks that are unlike any of the demonstrations. For instance, the agent never starts out holding an object in any of the training demonstrations, but it does so in about 50% of the testing task instances in the Burger environment. We observed that ViLa does not correctly put down the held object before attempting to pick up other objects, whereas pix2pred learns a predicate corresponding to the agent’s hands being empty and correctly selects this as a precondition for picking up any other object. We find that pix2pred only fails due to noise in VLM labeling: the VLM infrequently labels the initial state incorrectly such that a critical precondition is false, which prevents our planner from making a valid plan.
We also find that explicit optimization via hill-climbing leads to much better performance on testing tasks when compared to simply selecting predicates using a VLM (Q2). This can be seen from the fact that the VLM subselect baseline performs signifcantly worse than our main approach across all 4 task distributions. Qualitatively, we notice that this baseline tends to select far more predicates than necessary: in the burger domain, it selects predicates Cooked(?p1: patty) and Chopped(?l1: lettuce) (necessary for decision-making), but also extraneous predicates like Raw(?p1: patty), Available(?p1: patty), and Whole(?l1: lettuce) that the VLM labels quite inconsistently. This results in a complex set of operators that overfit strongly to the training tasks and do not generalize well. Additionally, we see that the “no subselect” baseline fails completely at all tasks, indicating that some subselection is critical; learning operators in terms of the more than 500 predicates in the proposal pool leads to extremely overfit operators.
We find that both feature-based and visual predicates are important in most of our tasks (Q3). The “no invent“ baseline fails completely in all the tasks because the initial predicates are insufficient. The “no visual.” baseline performs somewhat well in the Kitchen task, but fails in the Burger tasks where visual predicates, specifically predicates corresponding to Cooked and Chopped, are necessary, since the feature-based state does not contain sufficient information to classify these concepts. Interestingly, the predicate corresponding to whether the robot’s hand is empty can be classified more accurately in Burger tasks if features of the robot’s hand are used vs. attempting to recognize that concept visually. This leads to our approach doing better than the “no feat.” baseline in the Burger tasks, since our approach is not only limited to purely visual predicates and can invent whichever predicate has higher classification accuracy.
6 Limitations and Conclusion
@@
We proposed pix2pred, a method for inventing symbolic predicates that operate over raw images. We found that our approach, from just a handful of demonstrations (3-12), is able to invent semantically meaningful predicates that afford efficient planning and generalization to novel tasks across a range of domains and problems.
There are several noteworthy limitations of our present work. Firstly, VLM hallucinations adversely impact predicate proposal and atom labelling – especially with small amounts of demonstration data – which can inhibit our approach’s ability to learn useful predicates and operators. Secondly, the hill-climbing framework we use for predicate invention can be extremely slow, especially as the number of demonstrations and size of the initial predicate pool grows. Finally, our approach assumes input demonstrations are segmented in terms of a provided set of parameterized controllers with names that correspond to their function.
One interesting future direction is to expand the structure of predicates a VLM is allowed to produce, perhaps by allowing it to write arbitrarily-long text descriptions or even code to define new predicates (Liang et al. 2024). Another is to improve the underlying hill-climbing optimization procedure to be significantly more efficient and tolerant to noise, enabling pix2pred to scale to much larger domains that require potentially many tens or hundreds of predicates. Eventually we’d like to extend pix2pred to automatically segment demonstrations and learn controllers so that we can learn to plan directly from the simplest low-level demonstration data.
@@ Appendix A Appendix
A.1 Planner Implementation Details
1 Input: Task $\langle\mathcal{O},x_{0},g\rangle$, predicates $\Psi$, operators $\Omega$, samplers $\Sigma$.
2 $s_{0}\leftarrow\textsc{Abstract}(x_{0},\Psi)$
3 Call classical planner to generate an abstract plan: $\text{plan}^{\text{abs}}=\underline{\omega_{0}},\underline{\omega_{1}},...,\underline{\omega_{m}}\leftarrow\text{Planner}(s_{0},\mathcal{O},\Psi,\Omega,g)$
@@ 5 For $i=0,\dots,m$:
87  Use $\theta$ to fully ground controller $\underline{\mathcal{C}_{i}}$ into an action $a_{i}$. execute action $a_{i}$ and obtain the following state $x_{i+1}\leftarrow f(x_{i},a_{i})$
@@
Algorithm 1 shows the pseudocode for the planning and execution strategy we implement following recent work (Kumar et al. 2024b). Given an initial state $x_{0}$ for a task, we simply evaluate the classifiers of the given predicates $\Psi$ to convert the state into an abstract state $s_{0}$. We then call a classical planner with the task object set $\mathcal{O}$ and operators $\Omega$ to compute an abstract plan that achieves the goal (if one exists). We extract the ground controller sequence for this abstract plan, and then simply greedily execute each controller sequentially by calling the sampler associated with each operator.
We note that more sophisticated planning strategies are possible. In particular, provided a photorealistic simulation environment for the transition function $f$, we could leverage many task and motion planning (Garrett et al. 2021), such as bilevel planning (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022).
A.2 Operator Learning
We adapt the “cluster and intersect” operator learning strategy from previous work (Chitnis et al. 2022) to handle noise in predicate values inherent to our setting. Specifically, we learn a set of operators $\Omega$ from our demonstrations $\mathcal{D}$ and predicates $\Psi$ in four steps. Of these, the first three steps are largely taken directly from previous work: we introduce a modification to the third step, as well as the final step to handle noise.
@@ A.2 Operator Learning
    Partitioning: Each demonstration can be expressed as a sequence of transitions $\{(x,a,x^{\prime})\}$, with $x,x^{\prime}\in\mathcal{X}$ and $a\in\mathcal{A}$. Recall that each action $a$ is a controller with particular discrete and continuous arguments specified; let $\underline{\mathcal{C}}$ denote the corresponding controller with the same discrete object argument values, but continuous parameter values left unspecified. First, we use $\Psi$ to Abstract all states $x,x^{\prime}$ in the demonstrations $\mathcal{D}$, creating a dataset of transitions $\{(s,a,s^{\prime})\}$ with $s,s^{\prime}\in\mathcal{S}_{\Psi}$. Next, we partition these transitions via the following equivalence relation: $(s_{1},a_{1},s_{1}^{\prime})\equiv(s_{2},a_{2},s_{2}^{\prime})$ if the effects and controllers unify, that is, if there exists a mapping between the objects such that $\underline{\mathcal{C}_{1}}$, $(s_{1}-s_{1}^{\prime})$, and $(s_{1}^{\prime}-s_{1})$ are equivalent to $\underline{\mathcal{C}_{2}}$, $(s_{2}-s_{2}^{\prime})$, and $(s_{2}^{\prime}-s_{2})$ respectively. After this step, we have effectively ‘clustered’ all transitions in $\mathcal{D}$ together: we can associate each transition $\{(s,a,s^{\prime})\}$ with a particular equivalence class.
@@
    Arguments and Effects induction: For each equivalence class created in the previous step, we create $\overline{v}$ by selecting an arbitrary transition $(s,a,s^{\prime})$ and replacing each object that appears in the controller $\underline{\mathcal{C}}$ or effects with a variable of the same type. This further induces a substitution $\delta:\overline{v}\to\mathcal{O}$ for the objects $\mathcal{O}$ in this transition. Given this, the $E^{+}$, and $E^{-}$ can then be created by applying $\delta$ to $(s^{\prime}-s)$, and $(s-s^{\prime})$ respectively. By construction, for all other transitions $\tau$ in the same equivalence class, there exists an injective substitution $\delta_{\tau}$ under which the controller arguments and effects are equivalent to the newly created $E^{+}$, and $E^{-}$.
@@
    Precondition learning: The only remaining component required to turn each equivalence class into an operator is the operator preconditions. For this, we perform an intersection over all abstract states in each equivalence class (Bonet and Geffner 2019; Curtis et al. 2021). Recall that an abstract state is simply the collection of ground atoms that are ‘true’, thus taking an intersection amounts to finding the set of atoms that are always true across every initial state of every transition $(s,a,s^{\prime})$ in the equivalence class. However, since some of our predicate classifiers might be noisy they might not always hold. Thus, we take a ‘soft’ intersection: we take any atom as a precondition that is true across more than a specific percentage (set by a hyperparameter $h_{\text{pre\_frac}}$) of transitions in the equivalence class. More specifically, $P\leftarrow\bigcap_{\tau=(s,\cdot,\cdot)}\delta^{-1}_{\tau}(s)$ if $\frac{|\delta^{-1}_{\tau}(s)|}{|\mathcal{D}|}\geq h_{\text{pre\_frac}}$, where $\delta^{-1}_{\tau}(s)$ substitutes all occurrences of the objects in $s$ with the parameters in $\overline{v}$ following an inversion of $\delta_{\tau}$, and discards any atoms involving objects that are not in the image of $\delta_{\tau}$. $|\delta^{-1}_{\tau}(s)|$ denotes the number of transitions in $\mathcal{D}$ in which the lifted atom $|\delta^{-1}_{\tau}(s)|$ holds, and $|\mathcal{D}|$ denotes the total number of transitions in $\mathcal{D}$. In our experiments, we set $h_{\text{pre\_frac}}$ to 0.8.
@@
A.3 Sampler Learning
To enable execution, we must also learn samplers for proposing continuous controller parameters (Algorithm 1). Adapting prior work (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022), we train one sampler per operator, defined as:
$\sigma(x,o_{1},\dots,o_{k})=s_{\sigma}(x[o_{1}]\oplus\cdots\oplus x[o_{k}]),$
where $x[o]$ is the feature vector for $o$, $\oplus$ denotes concatenation, and $s_{\sigma}$ is a learned model. Treating this as supervised learning, we use operator-specific datasets $\tau_{\omega}$, where each of these datasets is composed of all transitions from the equivalence class used to induce this operator (Appendix A.2). Each transition $(s_{i},a_{i+1},s_{i+1})\in\tau_{\omega}$ maps operator arguments $\overline{v}$ to objects via $\delta:\overline{v}\to\mathcal{O}_{\tau}$. Using this mapping, the input for training is $x[\delta(v_{1})]\oplus\cdots\oplus x[\delta(v_{k})]$, where $(v_{1},\dots,v_{k})=\overline{v}$, and the output is the continuous parameter vector $\theta$, which are the continuous parameters used for $a_{i+1}$.
Each sampler is parameterized by two neural networks. The first predicts a Gaussian distribution over $\theta$, regressing its mean and covariance. The second network classifies $(x[o_{1}]\oplus\cdots\oplus x[o_{k}],\theta)$ as valid or invalid, enabling rejection sampling from the Gaussian. Negative examples are transitions outside $\tau_{\omega}$ but using the same controller.
A.4 More on Goals
The agent may sometimes invent visual predicates that can be interpreted as component of the task goal. For example, the goal of a task may be to make a burger with a cooked patty inside it. Here, the goal predicate would be an indicator for if this is achieved or not, perhaps called BurgerComplete(bottom_bun, patty, top_bun), in which case Cooked(patty) is an implicit component of that goal, in that BurgerComplete’s classifier implicitly checks that the patty is cooked. This is in contrast with a situation where the goal of a task is to fill a cup of water, and where the cup starts out upside down so that you can’t pour into it. Here, Upright(cup) may not be a component of HasWater(cup), in that the classifier for HasWater may not be explicitly checking if the cup is upright. In the former case, one could argue that there is no usefulness to inventing Cooked if we already have the goal predicate BurgerComplete, as we should be able to extract this concept from inside the goal predicate. But this may not be possible depending on how the classifier for the goal predicate was implemented – we may not have access to its internals. And even if we did, it may not be realistically possible to generate a dataset of burgers with and without cooked patties and label them with BurgerComplete’s classifier to obtain a dataset to train a classifier for Cooked. Furthermore, in our experimental setup, we would not necessarily need the classifier for BurgerComplete at all: at training time, any ground atoms associated with this predicate could be manually labeled in the demonstrations (easy to do when there are few goals per task), and at test time, because we execute one single plan without execution monitoring and replanning, we only need to check if the agent has solved the task once when the agent has completed its plan, which we could also do manually (analogously, the head chef at a restaurant does a manual check on the final burgers produced by the cooking staff).

### v4

@@ 5 Experiments
    No visual [42]: Ablation inventing only feature-based predicates ($\Psi^{\text{feat}}$), equivalent to.
  • $\bullet$
    No feature: Ablation inventing only visual predicates.
  • $\bullet$
    VLM feat. pred [16]: Baseline where a VLM is used to directly output predicates expressed as arbitrary code functions over $S^{\text{obj}}$ [13].
  • $\bullet$
    ViLa: Baseline where a VLM plans directly without abstractions; we test ViLa-pure (original version) and ViLa-fewshot (novel variant that adds demonstrations to the few-shot prompt).
Experimental Setup. In simulation, we vary object counts and locations across $5$ seeds, sample $10$ test tasks per seed, and allow one plan execution per task. In real-world domains, all demonstrations are collected from the point-of-view of a human, no object-centric state is available in the training data, and learned world models are evaluated on a robot. We compare our method to ViLA-fewshot using a fixed training set ($6$ demos for Cleanup, $10$ for Juice) and measure success as whether a correct plan is generated from a given initial state across $5$ generalization-focused test tasks. GPT-4o [35] is used as the VLM in simulation, and Gemini-2.0-flash [46] in the real world. All approaches use the same set of underlying skills in each domain. Specific details, including the precise prompts, skill implementations, and hyperparameters used can be found in Appendix A.6.
Task Approach New New More Novel Novel
obj. vis. obj. goal 1 goal 2
Cleanup Ours 3/3 3/3 3/3 3/3 3/3
ViLA-fewshot 3/3 3/3 2/3 0/3 3/3
Juice Ours 2/3 2/3 3/3 2/3 1/3
ViLA-fewshot 0/3 2/3 3/3 0/3 0/3
Table 1: Real-world generalization results. We evaluate on 5 different tasks testing different aspects of generalization in our 2 real world domains after training on the exact same dataset. Success out of 3 random initial states is reported. Additional details on task construction and evaluation are in Appendix A.6.
Results and Analysis. Figure 4 shows success rates in simulated domains. pix2pred outperforms prior approaches on 4 of 5 domains tested, often by a wide margin, and generalizes far better to complex tasks than ViLa few-shot (Q1). On the short-horizon kitchen task (2 actions), ViLa with demonstrations performs comparably to pix2pred, but its performance drops sharply on longer-horizon burger tasks (15+ actions). We find that ViLa tends to pattern-match training demonstrations and struggles with test tasks requiring true generalization—for example, handling initial states where the agent starts out holding an object (seen in $\sim$50% of Burger test tasks but never in training). ViLa fails to put down objects before picking up new ones, while pix2pred invents a HandEmpty predicate that enables the planner to handle such cases.
Our method also outperforms ViLA in real-world domains (Table 1). ViLA handles new backgrounds and object instances and quantities to some extent but struggles with novel goals requiring composition across demonstrations, especially in the complex Juice domain. In contrast, pix2pred generalizes across object instances, visual backgrounds, object counts, and novel goal compositions with just $6$ and $10$ human-collected demonstrations. Qualitatively, pix2pred invents few but meaningful predicates (e.g., OnFloor and IsEraser in Cleanup; JuiceMachineOpen, InsideWasteValveRegion, and NearJuiceValve in Juice). See Appendix A.8 for additional details.
We also find that explicit predicate subselection significantly improves test performance over direct VLM predicate selection (Q2). The VLM subselect baseline consistently underperforms pix2pred across all four task distributions. Qualitatively, VLM subselect picks many unnecessary predicates. In the Burger domain, for instance, it selects useful concepts like Cooked(?p1:patty) and Chopped(?l1:lettuce), but also inconsistent and redundant ones like Raw(?p1:patty), Available(?p1:patty), and Whole(?l1:lettuce). This bloats the operator set, causing overfitting and poor generalization. Finally, the “no subselect” baseline fails entirely, showing that subselection is critical. In most of our domains, the VLM proposes over $100$ distinct predicates; trying to learn operators over these directly leads to severe overfitting.
There are two possible causes of failure for our approach: (i) pix2pred learns a correct model, but fails to make a correct plan in test tasks because the VLM incorrectly classifies one or more predicates at test-time, and (ii) pix2pred fails to learn a model that generalizes to planning on test tasks (despite predicate values being “correct”). Upon manual inspection we found that all failures in our real-world environments, and the majority of failures in simulation environments, are due to (i).
6 Conclusion
@@
We proposed pix2pred, a method for inventing symbolic world models capable of enabling planning directly from input pixels. We found that our approach, from just a handful of demonstrations (3-12), is able to invent semantically meaningful predicates that afford efficient planning and aggressive generalization to novel object instances, numbers, task horizons, and goals across a range of problems.
7 Limitations
There are several important limitations of pix2pred in its current form. First, a key assumption of our approach is that, for each new task, all relevant objects in the scene are specified with unambiguous descriptors. This is currently done manually for the real-world environments and can be time-consuming. A second assumption is the full observability of objects in all states. These assumptions do not hold in all tasks of interest, especially when deploying in novel real-world environments. Future work could attempt to extract descriptors automatically from foundation models via a form of automatic prompt optimization [23], and extend the core ideas of pix2pred to partially observable settings.
A further limitation lies in the hill-climbing optimization procedure (Section 4.3) at the core of our model learning approach: it can be extremely slow — taking many seconds to evaluate the objective value of a single candidate predicate set — especially as the number of demonstrations and size of the initial predicate pool grow. We also found it to be quite sensitive to the choice of hyperparameters. Future work could develop more efficient and noise-tolerant optimization algorithms for predicate selection, as well as improved objectives for them. Relatedly, operator learning algorithms could be improved to better handle large input predicate sets and noisy predicates.
Finally, our approach assumes input demonstrations are segmented in terms of a provided set of parameterized skills with names that correspond to their function. These skills, as well as their underlying policies, are assumed to be provided ahead of time. Future work could integrate pix2pred with low-level skill-learning approaches, enabling learning from very low-level robot demonstrations or potentially even from large-scale internet video corpora.
Overall, we hope our work inspires future progress toward learning structured and interpretable world models that enable efficient inference in and strong generalization across complex robotic manipulation problems.
@@ Appendix A Appendix
A.1 Symbolic Operators
Given a particular state $s_{t}$, as well as predicates $\Psi$, we can compute the corresponding abstract state (denoted by $s^{\Psi}_{t}$) by selecting all the ground atoms that are “true” in $s_{t}$. Formally, $s^{\Psi}_{t}=\textsc{Abstract}(s_{t},\Psi)\triangleq\{\underline{\psi}:c_{\underline{\psi}}(s_{t})=\text{true},\forall\psi\in\Psi\}.$ Then, each operator $\omega$ is a tuple $\omega=\langle\overline{v},P,E^{+},E^{-},\mathcal{C}\rangle$. Here, $\overline{v}$ are typed variables representing the operator’s arguments, $P$, $E^{+}$ and $E^{-}$ are sets of predicates representing operator preconditions, add effects and delete effects respectively, and $\mathcal{C}$ is a controller associated with the operator. Note importantly that the discrete parameters of $\mathcal{C}$ must be a subset of $\overline{v}$. Specifying objects of the correct type as arguments induces a ground operator $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$. Ground operators define a (partial) abstract transition function: given $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$, if $\underline{P}\subseteq s$, then the successor abstract state $s^{\prime}$ is $(s\setminus\underline{E^{-}})\cup\underline{E^{+}}$ where $\setminus$ is set difference.
A ground operator specifies a corresponding ground controller $\underline{\mathcal{C}}$. To extract an action $a$ from this ground controller, we must specify values for the parameters $\Theta$ associated with controller $\mathcal{C}$. Following previous work [42, 26, 6], we leverage continuous parameter samplers to do this.
A.2 Planner Implementation Details
1Input: Task $\langle\mathcal{O},s_{0},g\rangle$, predicates $\Psi$, operators $\Omega$, samplers $\Sigma$.
2 $s^{\Psi}_{0}\leftarrow\textsc{Abstract}(s_{0},\Psi)$
3 Call classical planner to generate an abstract plan: $\text{plan}^{\text{abs}}=\underline{\omega_{0}},\underline{\omega_{1}},...,\underline{\omega_{m}}\leftarrow\text{Planner}(s^{\Psi}_{0},\mathcal{O},\Psi,\Omega,g)$
@@ 5 For $i=0,\dots,m$:
7  Use $\theta$ to fully ground controller $\underline{\mathcal{C}_{i}}$ into an action $a_{i}$.
8 execute action $a_{i}$ and obtain the following state $s_{i+1}\leftarrow f(s_{i},a_{i})$
@@
Algorithm 2 shows the pseudocode for the planning and execution strategy we implement following recent work [28]. Given an initial state $x_{0}$ for a task, we simply evaluate the classifiers of the Initial predicates $\Psi$ to convert the state into an abstract state $s_{0}$. We then call a classical planner with the task object set $\mathcal{O}$ and operators $\Omega$ to compute an abstract plan that achieves the goal (if one exists). We extract the ground controller sequence for this abstract plan, and then simply greedily execute each controller sequentially by calling the sampler associated with each operator.
We note that more sophisticated planning strategies are possible. In particular, provided a photorealistic simulation environment for the transition function $f$, we could leverage many task and motion planning [11], such as bilevel planning [42, 26, 6].
A.3 Operator Learning
We adapt the “cluster and intersect” operator learning strategy from previous work [6] to handle noise in predicate values inherent to our setting. Specifically, we learn a set of operators $\Omega$ from our demonstrations $\mathcal{D}$ and predicates $\Psi$ in four steps. Of these, the first three steps are largely taken directly from previous work: we introduce a modification to the third step, as well as the final step to handle noise.
@@ A.3 Operator Learning
    Partitioning: Each demonstration can be expressed as a sequence of transitions $\{(s,a,s^{\prime})\}$, with $s,s^{\prime}\in\mathcal{S}$ and $a\in\mathcal{A}$. Recall that each action $a$ is a controller with particular discrete and continuous arguments specified; let $\underline{\mathcal{C}}$ denote the corresponding controller with the same discrete object argument values, but continuous parameter values left unspecified. First, we use $\Psi$ to Abstract all states $s,s^{\prime}$ in the demonstrations $\mathcal{D}$, creating a dataset of transitions $\{(s^{\Psi},a,s^{\prime\Psi})\}$ with $s^{\Psi},s^{\prime\Psi}\in\mathcal{S}^{\Psi}$. Next, we partition these transitions via the following equivalence relation: $(s_{1}^{\Psi},a_{1},s_{1}^{\prime\Psi})\equiv(s_{2}^{\Psi},a_{2},s_{2}^{\prime\Psi})$ if the effects and controllers unify, that is, if there exists a mapping between the objects such that $\underline{\mathcal{C}_{1}}$, $(s_{1}^{\Psi}-s_{1}^{\prime\Psi})$, and $(s_{1}^{\prime\Psi}-s_{1}^{\Psi})$ are equivalent to $\underline{\mathcal{C}_{2}}$, $(s_{2}^{\Psi}-s_{2}^{\prime\Psi})$, and $(s_{2}^{\prime\Psi}-s_{2}^{\Psi})$ respectively. After this step, we have effectively ‘clustered’ all transitions in $\mathcal{D}$ together: we can associate each transition $\{(s^{\Psi},a,s^{\prime\Psi})\}$ with a particular equivalence class.
@@
    Arguments and Effects induction: For each equivalence class created in the previous step, we create $\overline{v}$ by selecting an arbitrary transition $(s^{\Psi},a,s^{\prime\Psi})$ and replacing each object that appears in the controller $\underline{\mathcal{C}}$ or effects with a variable of the same type. This further induces a substitution $\delta:\overline{v}\to\mathcal{O}$ for the objects $\mathcal{O}$ in this transition. Given this, the $E^{+}$, and $E^{-}$ can then be created by applying $\delta$ to $(s^{\prime\Psi}-s^{\Psi})$, and $(s^{\Psi}-s^{\prime\Psi})$ respectively. By construction, for all other transitions $\tau$ in the same equivalence class, there exists an injective substitution $\delta_{\tau}$ under which the controller arguments and effects are equivalent to the newly created $E^{+}$, and $E^{-}$.
@@
    Precondition learning: The only remaining component required to turn each equivalence class into an operator is the operator preconditions. For this, we perform an intersection over all abstract states in each equivalence class [4, 8]. Recall that an abstract state is simply the collection of ground atoms that are ‘true’, thus taking an intersection amounts to finding the set of atoms that are always true across every initial state of every transition $(s^{\Psi},a,s^{\prime\Psi})$ in the equivalence class. However, since some of our predicate classifiers might be noisy they might not always hold. Thus, we take a ‘soft’ intersection: we take any atom as a precondition that is true across more than a specific percentage (set by a hyperparameter $h_{\text{pre\_frac}}$) of transitions in the equivalence class. More specifically, $P\leftarrow\bigcap_{\tau=(s^{\Psi},\cdot,\cdot)}\delta^{-1}_{\tau}(s^{\Psi})$ if $\frac{|\delta^{-1}_{\tau}(s^{\Psi})|}{|\mathcal{D}|}\geq h_{\text{pre\_frac}}$, where $\delta^{-1}_{\tau}(s^{\Psi})$ substitutes all occurrences of the objects in $s$ with the parameters in $\overline{v}$ following an inversion of $\delta_{\tau}$, and discards any atoms involving objects that are not in the image of $\delta_{\tau}$. $|\delta^{-1}_{\tau}(s^{\Psi})|$ denotes the number of transitions in $\mathcal{D}$ in which the lifted atom $|\delta^{-1}_{\tau}(s^{\Psi})|$ holds, and $|\mathcal{D}|$ denotes the total number of transitions in $\mathcal{D}$. In our experiments, we set $h_{\text{pre\_frac}}$ to 0.8.
@@
A.4 Sampler Learning
To enable execution, we must also learn samplers for proposing continuous controller parameters (Algorithm 2). Adapting prior work [42, 26, 6], we train one sampler per operator, defined as:
$\sigma(s,o_{1},\dots,o_{k})=u_{\sigma}(s[o_{1}]\oplus\cdots\oplus s[o_{k}]),$
where $x[o]$ is the feature vector for $o$, $\oplus$ denotes concatenation, and $s_{\sigma}$ is a learned model. Treating this as supervised learning, we use operator-specific datasets $\tau_{\omega}$, where each of these datasets is composed of all transitions from the equivalence class used to induce this operator (Appendix A.3). Each transition $(s_{i}^{\Psi},a_{i+1},s_{i+1}^{\Psi})\in\tau_{\omega}$ maps operator arguments $\overline{v}$ to objects via $\delta:\overline{v}\to\mathcal{O}_{\tau}$. Using this mapping, the input for training is $s[\delta(v_{1})]\oplus\cdots\oplus s[\delta(v_{k})]$, where $(v_{1},\dots,v_{k})=\overline{v}$, and the output is the continuous parameter vector $\theta$, which are the continuous parameters used for $a_{i+1}$.
Each sampler is parameterized by two neural networks. The first predicts a Gaussian distribution over $\theta$, regressing its mean and covariance. The second network classifies $(s[o_{1}]\oplus\cdots\oplus s[o_{k}],\theta)$ as valid or invalid, enabling rejection sampling from the Gaussian. Negative examples are transitions outside $\tau_{\omega}$ but using the same controller.